# Ordnung muss sein

### v1

@@ Abstract.
Length category, ext-quiver, representation, partially ordered set
@@ 1. Introduction
A length category is an abelian category such that any chain of subobjects for a given object is finite and the isomorphism classes of objects form a set [6]. A basic example is given by the category of finite dimensional representations of a finite poset. Here, we work over a fixed field and the following result provides a characterisation of the length categories that arise in this way. [^1] Following the title of the paper, we may refer to conditions (1)–(3) as ‘Hausordnung’, cf. https://de.wikipedia.org/wiki/Hausmeister_Krause_-_Ordnung_muss_sein.
@@
Let $\mathbb{K}$ be a field. A $\mathbb{K}$-linear length category is equivalent to the category of finite dimensional representations of a finite poset if and only if the following holds:
@@
    The endomorphism ring of each simple object is isomorphic to $\mathbb{K}$.
@@
In this case the poset is given by the isomorphism classes of simple objects, with $S]\leq[T]$ if there is a path $S]\to\cdots\to[T]$ in the ext-quiver. In particular, the Hasse diagram of the poset identifies with the ext-quiver.
A fundamental discrete invariant of a length category is its ext-quiver. The vertices are given by the isomorphism classes of simple objects, and there is an arrow $S]\to[T]$ if $\operatorname{Ext}^{1}(T,S)\neq 0$. The above theorem describes the length categories that are completely determined by their ext-quiver.
There is a long tradition of studying representations of posets, and we refer to the monograph of Simson for a detailed account [9]. Intimately related is the study of incidence algebras and their modules. In fact, the question when a finite dimensional algebra is an incidence algebra has been investigated by several authors [1, 5, 8]. In particular, the relevance of distributive modules has been noticed, but the more general context of length categories seems to be new.
@@ 2. The proof
For any ring $A$ let $\operatorname{mod}A$ denote the category of finitely presented right $A$-modules. We write $\mathbb{K}P$ for the incidence algebra which is a $\mathbb{K}$-linear space with basis given by elements $a_{xy}$ for any pair $x\leq y$ in $P$ and multiplication induced by
$a_{xy}a_{vw}:=\begin{cases}a_{vy}&\text{if }w=x,\\ 0&\text{otherwise}.\end{cases}$
@@
We recall a few basic facts about modules over finite dimensional algebras. Let $A$ be a right artinian ring and fix a decomposition $1=\sum_{x\in P}e_{x}$ into pairwise orthogonal and primitive idempotents. We may assume that $A$ is basic. Thus the indecomposable projective modules are of the form $e_{x}A$ for some index $x$, and
@@
for any $A$-module $M$. Let $J=J(A)$ denote the Jacobson radical of $A$. Then the simple modules are of the form $S_{x}:=e_{x}(A/J)$ for some index $x$. Set $E_{x}:=\operatorname{End}_{A}(S_{x})$.
@@
We assume in addition that $A$ is a $\mathbb{K}$-algebra and that $E_{x}\cong\mathbb{K}$ for all $x$.
@@
Now fix an $A$-module $M$ such that $M:S_{x}]=1$ for all $x$, and we assume that every non-zero morphism $e_{x}A\to M$ is a monomorphism. Then $e_{x}A$ is multiplicity free for all $x$ and there is a $\mathbb{K}$-basis of $A$ that consists of elements $e_{xy}\in e_{y}Ae_{x}$. In particular, there is a partial order on the set $P$ given by
$x\leq y\quad:\iff\quad e_{y}Ae_{x}\neq 0,$
and for $x\leq y\leq z$ there is an element $\varepsilon_{xz}^{y}\in\mathbb{K}$ such that
$e_{yz}e_{xy}=e_{xz}\varepsilon_{xz}^{y}.$
Next we choose a $\mathbb{K}$-basis of $M$ that consists of elements $m_{x}\in Me_{x}$. Then there are non-zero elements $\mu_{xy}\in\mathbb{K}$ for all $x\leq y$ such that
$m_{y}e_{xy}=m_{x}\mu_{xy}.$
The equality
@@
gives
($\ast$) $\mu_{xz}\varepsilon_{xz}^{y}=\mu_{yz}\mu_{xy}.$
We are ready to compare $A$ with the incidence algebra $\mathbb{K}P$ that is given by the poset coming from the ext-quiver of $\operatorname{mod}A$.
@@
The $\mathbb{K}$-linear map $\phi\colon A\to\mathbb{K}P$ given by the assignment $e_{xy}\mapsto a_{xy}\mu_{xy}$ for any $x\leq y$ is an isomorphism of $\mathbb{K}$-algebras.
@@
The map is an isomorphism of vector spaces and we need to show that it is multiplicative. For $x\leq y\leq z$ we compute
$\phi(e_{yz}e_{xy})$ $=\phi(e_{yz}\varepsilon_{xz}^{y})$
$=a_{xz}\mu_{xz}\varepsilon_{xz}^{y}$
[…]
$=(a_{yz}a_{xy})\mu_{yz}\mu_{xy}$
$=(a_{yz}\mu_{yz})(a_{xy}\mu_{xy})$
$=\phi(e_{yz})\phi(e_{xy})$
and this yields the assertion. ∎
@@
Let $\mathcal{A}$ be a $\mathbb{K}$-linear length category. Suppose first that $\mathcal{A}$ is equivalent to $\operatorname{rep}(P,\mathbb{K})$ for some finite poset $P$. The simple representations identify with the elements of $P$. Specifically, for $x\in P$ the simple representation $S_{x}$ is given by $S_{x}(y)=\delta_{xy}\mathbb{K}$ for $y\in P$. In particular, it satisfies $\operatorname{End}(S_{x})\cong\mathbb{K}$. The ext-quiver identifies with the Hasse diagram of $P$. Thus there are no oriented cycles. Now consider the representation $M$ given by $M(x)=\mathbb{K}$ and $M(y)\to M(x)$ the identity map for all $x\leq y$. Then $M:S_{x}]=1$ for all $x$, and therefore the lattice of subobjects is distributive. The subrepresentation $M_{x}\subseteq M$ satisfying $M_{x}(y)\neq 0$ if and only if $y\leq x$ yields a projective cover of $S_{x}$. Thus every indecomposable projective representation embeds into $M$, and each object is a subquotient of $M^{r}$ for some $r\geq 1$.
Now suppose that conditions (1)–(3) hold for $\mathcal{A}$. It follows from [3, Proposition 2.14] that $\mathcal{A}$ is equivalent to the category $\operatorname{mod}A$ for some finite dimensional $\mathbb{K}$-algebra $A$. More precisely, conditions (2)–(3) imply that the category $\mathcal{A}$ admits a projective generator $Q$ such that $\operatorname{Hom}(Q,-)$ induces an equivalence $\mathcal{A}\xrightarrow{\raisebox{-1.20552pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}}\operatorname{mod}A$ for $A=\operatorname{End}(Q)$. Using Lemmas 2 and 5, the existence of a distributive object implies that Lemma 6 can be applied. Thus $A\cong\mathbb{K}P$ for the poset $P$ given by the isomorphism classes of simple $A$-modules, and the equivalence with $\operatorname{rep}(P,\mathbb{K})$ follows from Lemma 3. ∎
@@
The distributive object $M$ arising in condition (3) of Theorem 1 is not necessarily unique. In fact, such objects are parameterised by 22-cocycles $\mu=(\mu_{xy})$ in $H^{*}(\Sigma(P),\mathbb{K}^{\times})$, where $\Sigma(P)$ denotes the simplicial complex associated with $P$; see [7, 8].
@@
Condition (1) of Theorem 1 is needed. The category of finite length modules over the algebra $\mathbb{K}[\varepsilon]$ of duals numbers yields an example where (2) and (3) hold. Take for $M$ the indecomposable projective module.
The following example is interesting for several reasons, and I am grateful to Claus Ringel for suggesting it. The example shows that condition (2) of Theorem 1 is needed. Also, it yields infinitely many objects $M$ of length two such that all objects are a subquotient of a finite direct sum of copies of $M$.

### v2

@@ Abstract.
Length category, ext-quiver, centre, partially ordered set, linear representation, incidence algebra, distributive object
@@ 1. Introduction
A length category is an abelian category such that any chain of subobjects for a given object is finite and the isomorphism classes of objects form a set [8]. There are two basic invariants: the centre, which is the ring of natural transformations from the identity functor to itself, and the ext-quiver, for which the vertices are given by the isomorphism classes of simple objects, and there is an arrow $S]\to[T]$ if $\operatorname{Ext}^{1}(T,S)\neq 0$. The following result describes a class of length categories that are completely determined by these invariants. [^1] Following the title of the paper, we may refer to conditions (1)–(3) as ‘Hausordnung’, cf. https://de.wikipedia.org/wiki/Hausmeister_Krause_-_Ordnung_muss_sein.
For an abelian category $\mathcal{A}$ we write $Z(\mathcal{A})$ for its centre and say that an endomorphism of an object $X$ is central if it lies in the image of the canonical map $Z(\mathcal{A})\to\operatorname{End}(X)$. The category $\mathcal{A}$ is connected if 0 and $1$ are the only idempotents in $Z(\mathcal{A})$.
@@
A connected length category is equivalent to the category of finite dimensional representations of a finite poset over a field if and only if the following hold:
@@
    For each simple object all its endomorphisms are central.
@@
In this case the field equals the centre of the category and the poset is given by the isomorphism classes of simple objects, with $S]\leq[T]$ if there is a path from $[S]$ to $[T]$ in the ext-quiver.
There is a long tradition of studying linear representations of posets, and we refer to the monograph of Simson for a detailed account [11]. Intimately related is the study of incidence algebras and their modules. In fact, the question when a finite dimensional algebra is an incidence algebra has been investigated by several authors [2, 6, 10]. In particular, the relevance of distributive modules has been noticed, but the more general context of length categories seems to be new.
Categories of linear representations of finite posets over a field are nothing but categories of sheaves on finite $T_{0}$-spaces with values in the category of vector spaces; see Remark 14. This yields another perspective on the class of length categories arising in Theorem 1, which is reminiscent of Gabriel’s reconstruction of a noetherian scheme from its category of quasi-coherent sheaves [7, VI].
@@ 2. The proof
Module categories
We fix a length category $\mathcal{A}$ and recall from [4] a simple criterion for $\mathcal{A}$ to be a category of finite length modules. Let $\mathbb{K}$ be a commutative ring and suppose that $\mathcal{A}$ is $\mathbb{K}$-linear, so the morphism sets are $\mathbb{K}$-modules and the composition maps are bilinear. Then $\mathcal{A}$ is called hom-finite if $\operatorname{Hom}(X,Y)$ is a $\mathbb{K}$-module of finite length for each pair of objects $X,Y$.
For any ring $A$ let $\operatorname{mod}A$ denote the category of finitely presented right $A$-modules.
Proposition 3.
A hom-finite length category is equivalent to $\operatorname{mod}A$ for some Artin algebra $A$ if and only if there exists an object $M$ such that each object is a subquotient of a finite direct sum of copies of $M$.
Proof.
See Proposition 2.14 in [4]. ∎
Example 4.
Let $\mathcal{A}$ be a length category and view it as $\mathbb{K}$-linear category for $\mathbb{K}=Z(\mathcal{A})$. Suppose that for each simple object all its endomorphisms are central. Then $\mathcal{A}$ is hom-finite. More precisely, for each pair of objects $X,Y$ in $\mathcal{A}$ we have
$\ell_{\mathbb{K}}(\operatorname{Hom}(X,Y))\leq\ell(X)\cdot\ell(Y).$
@@
Let $\mathbb{K}P$ denote the incidence algebra which is a $\mathbb{K}$-linear space with basis given by elements $a_{xy}$ for any pair $x\leq y$ in $P$ and multiplication induced by
(1) $a_{xy}a_{vw}:=\delta_{wx}a_{vy}.$
@@
We recall a few basic facts about modules over finite dimensional algebras. Let $A$ be a right Artinian ring and fix a decomposition $1=\sum_{x\in P}e_{x}$ into pairwise orthogonal and primitive idempotents. We may assume that $A$ is basic and connected. Thus the indecomposable projective modules are of the form $e_{x}A$ for some index $x$, and
@@
for any $A$-module $M$. Let $J=J(A)$ denote the Jacobson radical of $A$. Then the simple modules are of the form $S_{x}:=e_{x}(A/J)$ for some index $x$. Set
$E_{x}:=\operatorname{End}_{A}(S_{x})\cong e_{x}(A/J)e_{x}.$
@@
Lemma 7.
Let $A$ be a right artinian ring such that the ext-quiver of $\operatorname{mod}A$ is connected and acyclic. Then the following are equivalent.
  1. (1)
    For each simple $A$-module all its endomorphisms are central.
  2. (2)
    The centre $Z(A)$ of $A$ is a field and isomorphic to the endomorphism ring of each simple $A$-module.
Proof.
One direction is clear. Thus assume that (1) holds. We may identify $Z(A)=Z(\operatorname{mod}A)$. For any index $x$ let $I_{x}$ denote the kernel of the canonical homomorphism $Z(A)\to E_{x}$. Then $Z(A)/I\xrightarrow{\raisebox{-1.20552pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}}E_{x}$ since we assume that all elements of $E_{x}$ are central. For each pair $x,y$ the action of $Z(A)$ on
$e_{y}Ae_{x}\cong\operatorname{Hom}_{A}(e_{x}A,e_{y}A)$
factors through $E_{x}$ and through $E_{y}$, since the ext-quiver is acyclic. Thus $I_{x}=I_{y}$ when $e_{y}Ae_{x}\neq 0$. Because $A$ is connected it follows that $I_{x}=I_{y}$ for all $x,y$. Finally, $I_{x}=0$ for all $x$ since $I_{x}$ annihilates all elements of $A$. ∎
From now on assume that for each simple $A$-module all its endomorphisms are central and set $\mathbb{K}:=Z(A)$.
@@
Now fix an $A$-module $M$ such that $M:S_{x}]=1$ for all $x$, and we assume that every non-zero morphism $e_{x}A\to M$ is a monomorphism. Then $e_{x}A$ is multiplicity free for all $x$ and there is a partial order on the set $P$ given by
$x\leq y\quad:\iff\quad e_{y}Ae_{x}\neq 0.$
Choose a $\mathbb{K}$-basis of $M$ that consists of elements $m_{x}\in Me_{x}$. This choice determines a $\mathbb{K}$-basis of $A$ consisting of elements $e_{xy}\in e_{y}Ae_{x}$ such that for all $x\leq y$
$m_{y}e_{xy}=m_{x}.$
For $x\leq y\leq z$ the equality
@@
implies
(2) $e_{yz}e_{xy}=e_{xz}.$
We are ready to compare $A$ with the incidence algebra $\mathbb{K}P$ that is given by the poset coming from the ext-quiver of $\operatorname{mod}A$. I am grateful to Dieter Vossieck for suggesting a simplification of this argument.
@@
The $\mathbb{K}$-linear map $A\to\mathbb{K}P$ given by the assignment $e_{xy}\mapsto a_{xy}$ for any $x\leq y$ is an isomorphism of $\mathbb{K}$-algebras.
@@
The map is an isomorphism of vector spaces; it is multiplicative because of the identities (1) and (2). ∎
@@
Let $\mathcal{A}$ be a length category. Suppose first that $\mathcal{A}$ is equivalent to $\operatorname{rep}(P,\mathbb{K})$ for some finite poset $P$ and a field $\mathbb{K}$. The simple representations identify with the elements of $P$. Specifically, for $x\in P$ the simple representation $S_{x}$ is given by $S_{x}(y)=\delta_{xy}\mathbb{K}$ for $y\in P$. In particular, it satisfies $\operatorname{End}(S_{x})\cong\mathbb{K}$. The ext-quiver identifies with the Hasse diagram of $P$. Thus there are no oriented cycles. Now consider the representation $M$ given by $M(x)=\mathbb{K}$ and $M(y)\to M(x)$ the identity map for all $x\leq y$. Then $M:S_{x}]=1$ for all $x$, and therefore the lattice of subobjects is distributive. The subrepresentation $M_{x}\subseteq M$ satisfying $M_{x}(y)\neq 0$ if and only if $y\leq x$ yields a projective cover of $S_{x}$. Thus every indecomposable projective representation embeds into $M$, and each object is a subquotient of $M^{r}$ for some $r\geq 1$.
Now suppose that conditions (1)–(3) hold for $\mathcal{A}$. It follows from Proposition 3 that $\mathcal{A}$ is equivalent to $\operatorname{mod}A$ for some Artin algebra $A$. More precisely, condition (2) implies that $\mathcal{A}$ is hom-finite; see Example 4. Then condition (3) implies that $\mathcal{A}$ admits a projective generator $Q$ such that $\operatorname{Hom}(Q,-)$ induces an equivalence $\mathcal{A}\xrightarrow{\raisebox{-1.20552pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}}\operatorname{mod}A$ for $A:=\operatorname{End}(Q)$. Next observe that $\mathbb{K}:=Z(\mathcal{A})$ is a field by Lemma 7. Here we need to assume that $\mathcal{A}$ is connected. Using Lemmas 2 and 8, the existence of a distributive object implies that Lemma 9 can be applied. Thus $A\cong\mathbb{K}P$ for the poset $P$ given by the isomorphism classes of simple $A$-modules, and the equivalence with $\operatorname{rep}(P,\mathbb{K})$ follows from Lemma 5. ∎
3. Remarks
We conclude with several comments on the main theorem.
@@ 3. Remarks
Consider a length category $\mathcal{A}=\operatorname{rep}(P,\mathbb{K})$ as in Theorem 1. The distributive object $M$ arising in condition (3) is not necessarily unique. In fact, such objects are parameterised by 22-cocycles $\mu=(\mu_{xy})$ in $H^{*}(\Sigma(P),\mathbb{K}^{\times})$, where $\Sigma(P)$ denotes the simplicial complex associated with $P$ and for $x\leq y$ the map $M(y)\to M(x)$ sends the basis element $m_{y}$ to $m_{x}\mu_{xy}$; see [9, 10].
@@
In Theorem 1 condition (1) is needed. Let $\mathcal{A}=\operatorname{mod}\mathbb{K}[\varepsilon]$ where $\mathbb{K}[\varepsilon]$ denotes the algebra of duals numbers over a field. This is not a category of poset representations, though conditions (2) and (3) hold. Take for $M$ the indecomposable projective module.
The following example is interesting for several reasons, and I am grateful to Claus Michael Ringel for suggesting it. The example shows that in Theorem 1 condition (2) is needed. Also, it yields infinitely many objects $M$ of length two such that all objects are a subquotient of a finite direct sum of copies of $M$.
@@
In Theorem 1 the assumption on the length category $\mathcal{A}$ to be connected can be removed. Then one obtains a finite product $\mathcal{A}_{1}\times\cdots\times\mathcal{A}_{n}$ of connected length categories and possibly different fields $Z(\mathcal{A}_{1}),\ldots,Z(\mathcal{A}_{n})$.
Remark 14.
Any partially ordered set $P$ carries the structure of a topological space, by taking as open subsets the downward closed subsets of $P$. This assignment identifies finite posets with finite $T_{0}$-spaces [1, §1]. Moreover, for any abelian category $\mathcal{C}$ the category of sheaves on $P$ with values in $\mathcal{C}$ identifies with the category of representations $P^{\mathrm{op}}\to\mathcal{C}$, by taking a sheaf $\mathscr{F}$ on $P$ to the representation
$x\longmapsto\mathscr{F}_{x}=\mathscr{F}(\downarrow\!x)\qquad(x\in P)$
where
$\downarrow\!x:=\{y\in P\mid y\leq x\}.$
Thus Theorem 1 characterises the categories of sheaves on finite $T_{0}$-spaces with values in the category of finite dimensional vector spaces over a field. One may compare this with Gabriel’s reconstruction of a noetherian scheme from its category of quasi-coherent sheaves [7, VI].
Remark 15.